# A shared-revenue Bertrand game

### v1

@@ D.2.2. Equilibria where $p\geq p_{r^{*}}$
More specifically, suppose $c_{s}\leq c_{s^{*}}$. We will show here the following claim: $p_{s^{*}}\leq p^{\dagger}$ for all $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$. Recalling our condition from (60), we have that
$p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,r}(p_{r^{*}})=\frac{1}{4}(1-c_{r})^{2}\leq\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)=\pi_{r,s}(p_{s^{*}},\alpha)$
In this case, we have
$\pi_{r,s}\left(p_{s^{*}},\alpha_{s^{*},r^{*}}\right)$ $=\frac{\alpha_{s^{*},r^{*}}}{4}\left(1-\left(\frac{c_{s}}{1-\alpha_{s^{*},r^{*}}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-\left(\frac{c_{s}}{\frac{c_{s}}{c_{r}}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-c_{r}^{2}\right)$
$\geq\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)(1-c_{r})(1+c_{r})$
$=\frac{(1-c_{r})^{2}}{4}=\pi_{r,r}(p_{r^{*}})$
Similarly, we can see that
$\pi_{r,s}\left(p_{s^{*}},\alpha^{\dagger}_{r}\right)$ $=\frac{\alpha^{\dagger}_{r}}{4}\left(1-\left(\frac{c_{s}}{1-\alpha^{\dagger}_{r}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(1-\left(\frac{c_{s}}{1-\frac{1-c_{r}}{1+c_{r}}}\right)^{2}\right)$
$\geq\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(1-\left(\frac{c_{s}}{1-\frac{c_{r}-c_{s}}{c_{r}}}\right)^{2}\right)$
$=\frac{(1-c_{r})^{2}}{4}=\pi_{r,r}(p_{r^{*}})$
Thus, to show the claim, all we need to show is that $\pi_{r,s}(p_{s^{*}},\alpha)$ is lower bounded by its value at the endpoints for $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$. We do this by finding the derivative:
$\frac{\partial\pi_{r,s}(p_{s^{*}},\alpha)}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}\left(\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)\right)$
$=\frac{1}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)-\frac{\alpha}{4}\left(\frac{2c_{s}^{2}}{(1-\alpha)^{3}}\right)$
$=\frac{1}{4}\left[1-\frac{c_{s}^{2}(1-\alpha)+2\alpha c_{s}^{2}}{(1-\alpha)^{3}}\right]$
(64) $=\frac{1}{4}\left[1-\frac{c_{s}^{2}(1+\alpha)}{(1-\alpha)^{3}}\right]$
Note that the $\tilde{\alpha}$ that maximizes $\pi_{r,s}(p_{s^{*}},\alpha)$ (which will appear in the future in Lemma 4 where we show that $\overline{\alpha}=\tilde{\alpha}$) satisfies the following –
$c_{s}^{2}=\frac{(1-\tilde{\alpha})^{3}}{1+\tilde{\alpha}}$
Because the function $f(x)=\frac{(1-x)^{3}}{1+x}$ is a bijection from (0,1][0,1)0101(0,1$\to$[0,1)( 0 , 1 ] → [ 0 , 1 ), this importantly implies that $\pi_{r,s}(p_{s^{*}},\alpha)$ has a unique maximum for $c_{s}^{2}\in[0,1)$. Now, we claim that $\pi_{r,s}$ is decreasing at $\alpha_{s^{*},r^{*}}$, which follows because –
00 0 $\leq\frac{\partial\pi_{r,s}(p_{s^{*}},\alpha)}{\partial\alpha}\biggr{|}_{\alpha=\alpha_{s^{*},r^{*}}}=\frac{1}{4}\left[1-\frac{c_{s}^{2}\left(\frac{2c_{r}-c_{s}}{c_{r}}\right)}{\left(\frac{c_{s}}{c_{r}}\right)^{3}}\right]$
00 0 $\leq 1-\frac{c_{r}^{2}(2c_{r}-c_{s})}{c_{s}}$
$\frac{c_{s}}{c_{r}^{2}}$ $\leq 2c_{r}-c_{s}$
$\alpha_{s^{*}}=1-2c_{r}+c_{s}$ $\leq\frac{c_{r}-c_{s}}{c_{r}^{2}}=\frac{\alpha_{s^{*},r^{*}}}{c_{r}}$
Which always holds because, as we showed in (54), when $c_{s}\leq c_{s^{*}}$ we have $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}\leq\frac{\alpha_{s^{*},r^{*}}}{c_{r}}$ always. This is sufficient to conclude the proof of the claim. Given the claim, we can now conclude that when $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$, the region of equilibria we have found is $p\in\left[p_{s_{\text{ind}}},p_{s^{*}}\right]$, which is the same as we found before. Note that $p^{\dagger}$ is well defined for all $\alpha\geq(1-c_{r})^{2}$, which is implied by $\alpha\geq\alpha^{\dagger}_{r}$:
(65) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\geq(1-c_{r})^{2}$
11 11 $\geq(1-c_{r})(1+c_{r})$
$c_{r}^{2}$ $\geq 0$
Reconciling equilibria with case (ii) where $c_{s}\geq c_{s^{*}}$.
Switching the direction of the inequalities in the proof from the previous subsection yields the analogous result that if $c_{s}\geq c_{s^{*}}$, $p^{\dagger}\leq p_{s^{*}}$ for all $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right]$. We previously found that for all $\alpha\leq\alpha^{\dagger}_{r}$, the equilibria are $p\in\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]$. However, by the same argument as in (60) we have that
$\frac{1}{2}\leq p^{\dagger}\leq p_{r_{\text{ind}}}\iff\pi_{r,s}(p^{\dagger})\geq\pi_{r,s}(p_{r_{\text{ind}}})\iff\pi_{r,r}(p_{r^{*}})\geq\pi_{r,r}(p_{r_{\text{ind}}})$
@@
which always holds, thus we have found no new equilibria in this region of parameter space.
@@
(66) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha_{s^{*}}\\ \left[p_{s_{\text{ind}}},p_{s^{*}}\right]&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ \left[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]&\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\\ \end{cases}$
@@
(67) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha^{\dagger}_{r}\\ \left[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}\\ \end{cases}$
@@ D.3. Equilibria where $p_{r}<p_{s}$
Condition IV is always satisfied; because for all $\varepsilon>0$, $p_{s}+\varepsilon>p_{s}>p_{r}$, the independent seller achieves a payoff of 00 both at price $p_{s}$ and price $p_{s}+\varepsilon$. Together, conditions I and II imply that the retailer must set price $p_{r^{*}}$ so as to maximize their payoff function. Otherwise, if they are not at the optimal price, they would be able to increase their payoff by changing their price.
Furthermore, condition II requires that it is not the retailer would not achieve a higher payoff by deviating to some price $p_{r}>p_{s}$ so as to let the independent seller fulfill demand in equilibrium. Formally, we need that
@@
However, this is exactly the equation we solved in (58), so we have that $p_{s}$ must be greater than $p^{\dagger}$ (recall from (59) that the smaller root is always less than $p_{r^{*}}$). Note that we also must have that $p_{s}\geq p_{r^{*}}$, and in (63) we showed that $p_{r^{*}}\leq p^{\dagger}$ is equivalent to $\alpha\geq\alpha^{\dagger}_{r}$.
@@
The maximum of the independent seller’s payoff on the interval $[0,p_{r^{*}}]$ is attained at
@@
because $\pi_{s}(\cdot,p_{s}^{\prime})$ is maximized at $p_{s^{*}}$. Consider the second case when $p_{r^{*}}>p_{s^{*}}$. Recall from (9) that
$\pi_{s,s}(p_{s^{*}})=\begin{cases}\frac{1-\alpha}{4}\left(1-\frac{c_{s}}{1-\alpha}\right)^{2}&\alpha<1-c_{s}\\ 0&\alpha\geq 1-c_{s}\end{cases}$
Thus, we find immediately that when $p_{r^{*}}>p_{s^{*}}$, condition III never holds if $\alpha<1-c_{s}$ and always holds if $\alpha\geq 1-c_{s}$. However, if $\alpha\geq 1-c_{s}$, we have that $p_{s^{*}}\geq 1>p_{r^{*}}$, which is a contradiction. Thus there are no equilibria where $p_{r^{*}}>p_{s^{*}}$.
On the other hand, $\arg\max_{p_{s}^{\prime}\in[0,p_{r^{*}}]}\left[\pi_{s}(p_{r^{*}},p_{s}^{\prime})\right]=p_{r^{*}}$ if $p_{r^{*}}\leq p_{s^{*}}$, which we showed in (52) is equivalent to $\alpha\geq\alpha_{s^{*},r^{*}}$. With this, condition III is satisfied in this case if
$\pi_{s,s}(p_{r^{*}})$ $\leq 0$
$\left((1-\alpha)\left(\frac{1+c_{r}}{2}\right)-c_{s}\right)\left(1-\frac{1+c_{r}}{2}\right)$ $\leq 0$
$\left(\frac{(1-\alpha)(1+c_{r})-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$ $\leq 0$
$(1-\alpha)(1+c_{r})$ $\leq 2c_{s}$
(68) $\frac{1+c_{r}}{2}$ $\leq\frac{c_{s}}{1-\alpha}$
(69) $\alpha$ $\geq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
As a remark, see that (68) can be rewritten as $p_{r^{*}}\leq p_{s_{\text{ind}}}$, which is obvious in hindsight because $p_{s_{\text{ind}}}$ is by construction the price below which the independent seller makes nonpositive payoff. We showed in (51) that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$, so $\alpha\geq\alpha_{r^{*}}$ is sufficient to imply that $p_{r^{*}}\leq p^{\dagger}$ per (63). Additionally, recall from (53) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$.
In summary, we have a continuum of equilibria $(p_{r^{*}},p_{s}\geq p^{\dagger})$ whenever $\alpha\geq\alpha_{r^{*}}$.
@@ D.4. Equilibria where $p_{s}<p_{r}$
In this case, condition II is always satisfied; the retailer does not fulfill any demand either way as, for all $\varepsilon>0$, $p_{r}+\varepsilon>p_{r}>p_{s}$. As in the previous case, conditions III and IV together imply that the independent seller must set price $p_{s^{*}}$. Condition IV requires that the independent seller does not prefer to let the retailer fulfill demand in equilibrium, in which case they would earn a payoff of 00. To satisfy this, we need only that the independent seller sets a price $p_{s}\geq c_{s}$, which always holds for price $p_{s^{*}}$.
@@
Once again, there are two cases for the choice of $p_{r}^{\prime}$ that maximizes the retailer’s payoff function on the interval $[0,p_{s^{*}}]$
@@
Consider the second case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{r^{*}}$. From (52), recall that $p_{s^{*}}>p_{r^{*}}\iff\alpha>\alpha_{s^{*},r^{*}}$. Recalling our analysis from (60), condition I then simplifies to
@@
Now consider the first case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{s^{*}}$. From (52), recall that $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$. On the other hand, we have that condition I becomes
$\pi_{r,r}(p_{s^{*}})$ $\leq\pi_{r,s}(p_{s^{*}})$
$(1-p_{s^{*}})(p_{s^{*}}-c_{r})$ $\leq\alpha p_{s^{*}}(1-p_{s^{*}})$
$(1-\alpha)p_{s^{*}}$ $\leq c_{r}$
$\frac{1-\alpha+c_{s}}{2}$ $\leq c_{r}$
$\alpha$ $\geq 1-2c_{r}+c_{s}=\alpha_{s^{*}}$
Thus, we have a continuum of equilibria $(p_{r}>p_{s^{*}},p_{s^{*}})$ whenever $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$. Note that it’s not always the case that $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}$, but when this holds, such equilibria exist. Additionally, it’s clear that $\alpha_{s^{*},r^{*}}=1-\frac{c_{s}}{c_{r}}<1-c_{s}$ always.
@@ E.1. Admissibility
Thus, in any region of parameter space, the only admissible strategies for the independent seller are $p_{s}\in(p_{s_{\text{ind}}},p_{s^{*}})$, except for a technical edge case when $p_{s_{\text{ind}}}\geq p_{s^{*}}\geq 1$ which we will address later.
A similar result holds for the retailer, except it’s not always the case that $p_{r_{\text{ind}}}\leq p_{r^{*}}$ (which occurs precisely when $\alpha\leq\alpha^{\dagger}_{r}$ per (49)). We claim that $p_{r}=\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}$ dominates all $p_{r}^{\prime}<p_{r}$. Indeed, for $p_{s}\leq p_{r}^{\prime}$, the retailer’s payoff is the same. For $p_{s}\geq p_{r}$, the retailer’s payoff is
@@
The arguments of the previous paragraph did not depend on the relative ordering of $p^{\dagger}$ and $p_{s^{*}}$, and we know that $p_{r_{\text{ind}}}\geq p^{\dagger}\geq p_{r^{*}}$ per (63) because $\alpha\geq\alpha^{\dagger}_{r}$. Thus the admissible equilibria when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha_{r^{*}}\right]$ and $p^{\dagger}\leq p_{s^{*}}$ are simply the shared prices $p\in\left[p_{r^{*}},p^{\dagger}\right]$.
@@
However, when $\alpha\geq 1-c_{s}\geq\alpha^{\dagger}_{s}$, we have that $p_{s_{\text{ind}}}\geq p_{s^{*}}\geq 1$ and the Nash equilibrium is again given by $(p_{r}=p_{r^{*}},p_{s}\geq p^{\dagger})$. In this limit, setting $p_{s}=1$ weakly dominates setting any $p_{s}^{\prime}<1$ because $p_{s_{\text{ind}}}\geq 1>p_{s}^{\prime}$ which implies that the independent seller would achieve negative payoff for each unit sold at price $p_{s}^{\prime}$ and would prefer to achieve 0 payoff by setting $p_{s}=1$. Thus, the only admissible equilibrium when $\alpha\geq 1-c_{s}$ is $(p_{r}=p_{r^{*}},p_{s}=1)$. Note that this edge case is technical and unimportant, as no matter what price the independent seller sets, the result remains the same – when $\alpha$ is high, there is an equilibrium where the retailer fulfills demand.
@@
(70) $(p_{r},p_{s})\in\begin{cases}(p_{s},[p_{s_{\text{ind}}},p^{\dagger}])\cup(p_{r^{*}},[p^{\dagger},p_{s^{*}}])&\left[\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land[p^{\dagger}\leq p_{s^{*}}]\\ (p_{s},[p_{s_{\text{ind}}},p_{s^{*}}])&\left[\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land[p_{s^{*}}\leq p^{\dagger}]\\ (p_{r^{*}},[p_{s_{\text{ind}}},p_{s^{*}}])&\alpha^{\dagger}_{s}\leq\alpha\leq 1-c_{s}\\ (p_{r^{*}},1)&1-c_{s}\leq\alpha\\ \end{cases}$
@@
(71) $(p_{r},p_{s})\in\begin{cases}(p_{r_{\text{ind}}},p_{r_{\text{ind}}})&\alpha\leq 1-2c_{r}+c_{s}\\ ([p_{r_{\text{ind}}},p_{r^{*}}],p_{s^{*}})&\alpha_{s^{*}}\leq\alpha\leq\alpha^{\dagger}_{r}\\ ([p_{r^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ (p_{s},\left[p_{r^{*}},p_{s^{*}}\right])\cup([p_{s^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{r^{*}},p^{\dagger}\right])&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ \end{cases}$
@@
(72) $(p_{r},p_{s})\in\begin{cases}(p_{r_{\text{ind}}},p_{r_{\text{ind}}})&\alpha\leq\alpha^{\dagger}_{r}\\ (p_{s},\left[p_{r^{*}},p_{s^{*}}\right])\cup([p_{s^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{r^{*}},p^{\dagger}\right])&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ \end{cases}$

### v2

@@ D.2.2. Equilibria where $p\geq p_{r^{*}}$
Suppose $c_{s}\geq c_{s^{*}}$ and $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right]$. Then, we claim this implies $p^{\dagger}<p_{s^{*}}$. Suppose for contradiction $p_{s^{*}}\leq p^{\dagger}$. Then, since $\alpha<\alpha^{\dagger}_{r}\leq\alpha_{s^{*}}$, we have $p_{r_{\text{ind}}}\leq p_{s^{*}}<p^{\dagger}$, which implies
$p^{\dagger}>p_{r_{\text{ind}}}\implies\pi_{r,r}(p^{\dagger})>\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$
contradicting the optimality of $p_{r^{*}}$. Now recall that earlier, we found that for all $\alpha\leq\alpha^{\dagger}_{r}$, the equilibria are $p\in\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]$. However, we have that
$\tilde{p}\leq p^{\dagger}\leq p_{r_{\text{ind}}}\iff\pi_{r,s}(p^{\dagger})\geq\pi_{r,s}(p_{r_{\text{ind}}})\iff\pi_{r,r}(p_{r^{*}})\geq\pi_{r,r}(p_{r_{\text{ind}}})$
@@
which always holds, thus we have found no new equilibria in this region of parameter space since $p_{s_{\text{ind}}},p_{r_{\text{ind}}}]\supseteq[p_{r^{*}},p^{\dagger}]$.
Equilibria when $c_{s}\leq c_{s^{*}}$.
Now suppose $c_{s}\leq c_{s^{*}}$ and $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$, we claim this implies $p_{s^{*}}\leq p^{\dagger}$. Suppose for contradiction $p^{\dagger}<p_{s^{*}}$. Since $\alpha\geq\alpha^{\dagger}_{r}$, we have $p_{r^{*}}\leq p^{\dagger}<p_{s^{*}}$, but this contradicts the fact that $\alpha\leq\alpha_{s^{*},r^{*}}\iff p_{s^{*}}\leq p_{r^{*}}$. Thus, in this region of parameter space, we have found equilibria $p\in[p_{r^{*}},p_{s^{*}}]$, but note that $\alpha\leq\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$ which implies $p_{s_{\text{ind}}}\leq p_{r^{*}}$. Therefore we have found no new equilibria once again in this region once again because before the equilibria we found were $p\in[p_{s_{\text{ind}}},p_{r^{*}}]\supseteq[p_{s^{*}},p_{r^{*}}]$.
@@
[…]
@@
[…]
Note that $\alpha\geq\alpha^{\dagger}_{r}$ whenever $p^{\dagger}$ is a possible equilibrium, which is sufficient for $p^{\dagger}$ to be well-defined per Appendix C.3.
@@ D.3. Equilibria where $p_{r}<p_{s}$
Condition IV is always satisfied because for all $\varepsilon>0$, $p_{s}+\varepsilon>p_{s}>p_{r}$, the independent seller achieves a payoff of 0 both at price $p_{s}$ and price $p_{s}+\varepsilon$. Together, conditions I and II imply that the retailer must set price $p_{r^{*}}$ so as to maximize their payoff function. Otherwise, if they are not at the optimal price, they would be able to increase their payoff by changing their price.
Furthermore, condition II requires that the retailer would not achieve a higher payoff by deviating to some price $p_{r}>p_{s}$ so as to let the independent seller fulfill demand in equilibrium. Formally, we need that
@@
However, by construction this is equivalent to $p_{s}\geq p^{\dagger}$ (note that the smaller root is less than $\tilde{p}$ and therefore $p_{r^{*}}$ by Lemma 2). Note that we also must have that $p_{s}\geq p_{r^{*}}$, and we have $p_{r^{*}}\leq p^{\dagger}\iff\alpha\geq\alpha^{\dagger}_{r}$.
@@
The maximum of the independent seller’s payoff on the interval $0,p_{r^{*}}]$ is attained at
@@
because $\pi_{s}(\cdot,p_{s}^{\prime})$ is maximized at $p_{s^{*}}$. Consider the second case when $p_{r^{*}}>p_{s^{*}}>p_{s_{\text{ind}}}$. Then, there always exists some price in $(p_{s_{\text{ind}}},p_{r^{*}})$ where the independent seller can make positive payoff so they would prefer to undercut the independent seller. Thus there are no equilibria where $p_{r^{*}}>p_{s^{*}}$.
On the other hand, $\arg\max_{p_{s}^{\prime}\in[0,p_{r^{*}}]}\left[\pi_{s}(p_{r^{*}},p_{s}^{\prime})\right]=p_{r^{*}}$ if $p_{r^{*}}\leq p_{s^{*}}$, which is equivalent to $\alpha\geq\alpha_{s^{*},r^{*}}$. In order for this to be an equilibrium, we need further that $p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}$ so that the independent seller does not prefer to undercut the retailer.
We showed in (41) that $\alpha^{\dagger}_{r}<\alpha_{r^{*}}$, so $\alpha\geq\alpha_{r^{*}}$ is sufficient to imply that $p_{r^{*}}<p^{\dagger}$. Additionally, recall from (40) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$. In summary, we have a continuum of equilibria $(p_{r^{*}},p_{s}\geq p^{\dagger})$ whenever $\alpha\geq\alpha_{r^{*}}$.
@@ D.4. Equilibria where $p_{s}<p_{r}$
In this case, condition II is always satisfied; the retailer does not fulfill any demand either way as, for all $\varepsilon>0$, $p_{r}+\varepsilon>p_{r}>p_{s}$. As in the previous case, conditions III and IV together imply that the independent seller must set price $p_{s^{*}}$. Condition IV also requires that the independent seller does not prefer to let the retailer fulfill demand in equilibrium, in which case they would earn a payoff of 0. To satisfy this, we need only that the independent seller sets a price $p_{s}\geq c_{s}$, which always holds for price $p_{s^{*}}$.
@@
Once again, there are two cases for the choice of $p_{r}^{\prime}$ that maximizes the retailer’s payoff function on the interval $0,p_{s^{*}}]$
@@
Consider the second case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{r^{*}}$. Recall that $p_{s^{*}}>p_{r^{*}}\iff\alpha>\alpha_{s^{*},r^{*}}$. Recalling Lemma 2, condition I then simplifies to
@@
Now consider the first case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{s^{*}}$. Recall that by construction, $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$. On the other hand, we have that condition I becomes
$\pi_{r,r}(p_{s^{*}})\leq\pi_{r,s}(p_{s^{*}})\iff p_{s^{*}}\leq p_{r_{\text{ind}}}\iff\alpha\geq\alpha_{s^{*}}$
Thus, we have a continuum of equilibria $(p_{r}>p_{s^{*}},p_{s^{*}})$ whenever $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$. Note that it’s not always the case that $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}$, but when this holds, such equilibria exist.
@@ E.1. Admissibility
Thus, in any region of parameter space, the only admissible strategies for the independent seller are $p_{s}\in(p_{s_{\text{ind}}},p_{s^{*}})$, except for a technical edge case when $\alpha\geq 1-c_{s}\implies p_{s_{\text{ind}}}\geq 1$ which we will address later.
A similar result holds for the retailer, except it’s not always the case that $p_{r_{\text{ind}}}\leq p_{r^{*}}$ (which occurs precisely when $\alpha\leq\alpha^{\dagger}_{r}$ by construction). We claim that $p_{r}=\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}$ dominates all $p_{r}^{\prime}<p_{r}$. Indeed, for $p_{s}\leq p_{r}^{\prime}$, the retailer’s payoff is the same. For $p_{s}\geq p_{r}$, the retailer’s payoff is
@@
The arguments of the previous paragraph did not depend on the relative ordering of $p^{\dagger}$ and $p_{s^{*}}$, and we know that $p_{r_{\text{ind}}}\geq p^{\dagger}\geq p_{r^{*}}$ because $\alpha\geq\alpha^{\dagger}_{r}$. Thus the admissible equilibria when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha_{r^{*}}\right]$ and $p^{\dagger}\leq p_{s^{*}}$ are simply the shared prices $p\in\left[p_{r^{*}},p^{\dagger}\right]$.
@@
However, when $\alpha\geq 1-c_{s}\geq\alpha^{\dagger}_{s}$, we have that $p_{s_{\text{ind}}}\geq 1$ and the Nash equilibrium is again given by $(p_{r}=p_{r^{*}},p_{s}\geq p^{\dagger})$. In this limit, setting $p_{s}=1$ weakly dominates setting any $p_{s}^{\prime}<1$ because $p_{s_{\text{ind}}}\geq 1>p_{s}^{\prime}$ which implies that the independent seller would achieve negative payoff for each unit sold at price $p_{s}^{\prime}$ and would prefer to achieve 0 payoff by setting $p_{s}=1$. Thus, the only admissible equilibrium when $\alpha\geq 1-c_{s}$ is $(p_{r}=p_{r^{*}},p_{s}=1)$. Note that this edge case is technical and unimportant, as no matter what price the independent seller sets, the result remains the same – when $\alpha$ is high, there is an equilibrium where the retailer fulfills demand.
@@
[…]
@@
[…]
@@
[…]